OpenStudy (seattle12345):

Can someone help with this? Please:Write 5 cos^2 (20 a) + 5 sin^2(20 a) as one trigonometric function.

OpenStudy (mathstudent55):

Do you know the identity: \(\\sin^2 \theta + \cos^2 \theta = 1\) Factor out a 5. What do you get?

OpenStudy (seattle12345):

\[5(\cos ^{2}\left( 20 a\right)+\sin ^{2}(20a))\]?

OpenStudy (ranga):

The one inside the parenthesis equals 1 because of the identity listed in the first reply.

OpenStudy (seattle12345):

oh ok I see it.

OpenStudy (mathstudent55):

Now apply the identity to what is inside the parentheses. \(5 \cos^2 20 a + 5 \sin^2 20 a\) \(=5 \color{red}{( \cos^2 20 a + \sin^2 20 a )} \) \(=5 \cdot \color{red}1 \) \( = 5\)

OpenStudy (seattle12345):

Thank you! I guess (20a) was confusing me a bit, but I see the identity

OpenStudy (mathstudent55):

Right. The sine squared of anything plus the cosine squared of the same thing equals one, no matter what the thing is. In your case, the thing was 20a.